a)If 7.31 moles of Ethane(C2H6) undergoes combustion according to the unbalanced equation C2H6+O2->CO2+H20, how much oxygen is required? Answer in units of mol.

b)How much CO2 is formed? Answer in units of mol.

c)How much H2O is formed? answer in units of mol.
1
Expert's answer
2013-12-21T08:19:44-0500
Write down the balanced equation of the process:

2C2H6 + 7O2 -> 4CO2 + 6H2O

According to the equation,
n(O2) = 7/2*n(C2H6) = 25.6 mol
n(CO2) = 4/2*n(C2H6) = 14.6 mol
n(H2O) = 6/2*n(C2H6) = 21.9 mol